A natural monopolist faces the following demand curve: P = 409 - 2Q, its total cost is given by: TC = 12800 + 9Q (marginal cost is the slope of total cost).

(a) If the government regulates the monopolist to charge a socially optimal price, what price will it charge and how many units will it sell? How much are the profit, consumer surplus and producer surplus?

(b) If it is not a regulated monopolist, what is its profit maximizing price and quantity (assuming single price monopolist)? How much are the profit, consumer surplus, producer surplus and dead weight loss?

(c) If the regulated monopolist is allowed to charge a fair return price (which is $89), how much are the consumer surplus, producer surplus and dead weight loss?
Answer #1

aws h o I socially 409 - 20, TC= 19800 +90 optimal price = MC = atclaa= $9, a) ET then from dd aowe P = 409-20 Q= 409-1) 12 =elated Monopolist, P = 89, where dd awowe auts AC, - then = (409-89/2 = 160 • CS = 2x 160 al 409-9) = * X 160X (409-89) = (16